# Tunable Entanglement in Cavity-Magnon Optomechanics

### v2

@@ II.3 Covariance matrix
where $\mathcal{D}=\operatorname{diag}[\kappa_{1}(2\bar{n}_{1}+1),\kappa_{1}(2\bar{n}_{1}+1),\kappa_{2}(2\bar{n}_{2}+1),\kappa_{2}(2\bar{n}_{2}+1),\kappa_{a}(2\bar{n}_{a}+1),\kappa_{a}(2\bar{n}_{a}+1),0,\gamma_{b}(2\bar{n}_{b}+1)]$ is defined by $\langle\bar{n}_{l}(t)\bar{n}_{k}(t^{\prime})$ $+\bar{n}_{k}(t^{\prime})\bar{n}_{l}(t)\rangle=2\mathcal{D}_{lk}\delta(t-t^{\prime})$.
@@ III Optomechanical entanglement
In this part, we study the magnon-mediated optomechanical entanglement. To be more clear, we first re-examine the optomechanical entanglement without magnons, i.e., $g_{1}=0$ and $g_{2}=0$. Then we investigate the magnon-mediated optomechanical entanglement by taking a single sphere for example. Without loss of generaity, we assume $g_{1}=g_{m}\neq 0$ and $g_{2}=0$.
@@ III.1 Without magnons
Figure 2: (a) The optomechanical entanglement ($E_{N}^{ab}$) vs the normalized cavity detuning $\tilde{\Delta}_{a}/\omega_{b}$ with different optomechanical coupling strengths $G/\kappa_{a}=1,3,5$ at the bath temperature $T=20$ mK. (b) The LN $E_{N}^{ab}$ vs the bath temperature $T$ with the moderate optomechanical coupling strength $G/\kappa_{a}=4$ and the frequency detuning $\tilde{\Delta}_{a}=0.9\omega_{b}$. Other parameters are chosen as $\omega_{a}/2\pi=10$ GHz, $\omega_{b}/2\pi=10$ MHz, $\kappa_{a}/2\pi=1$ MHz, $\gamma_{b}/2\pi=100$ Hz.
@@
When no spheres are placed in the cavity (i.e., $g_{1}=g_{2}=0$), the proposed cavity-magnon optomechanics reduces to a standard optomechanical configuration comprising solely a cavity and a mechanical resonator. In this setup, optomechanical entanglement can be analyzed by plotting LN ($E_{N}^{ab}$) vs the normalized cavity detuning $\tilde{\Delta}_{a}/\omega_{b}$ with different linearized coupling strengths $G$ in Fig. 2(a). Here, we employ experimentally accessible parametersZhang et al. (2016); Kleckner et al. (2006): $\omega_{a}/2\pi$ = 10 GHz, $\omega_{b}/2\pi$ = 10 MHz, $\kappa_{a}/2\pi=1$ MHz, $\gamma_{b}/2\pi=10^{2}$ Hz, and $T=20$ mK. Evidently, the optomechanical entanglement increases first and then decreases with the cavity detuning $\tilde{\Delta}_{a}/\omega_{b}$. Around $\tilde{\Delta}_{a}\approx\omega_{b}$, the optimal optomechanical entanglement is observed. By gradually increasing the optomechanical coupling, the optomechanical entanglement can be significantly enhanced (as depicted by the three curves). The generated optomechanical entanglement arises from both the two-mode squeezing effect ($a^{\dagger}b^{\dagger}+ab$, where $q\propto b+b^{\dagger}$) and the beam-splitter effect ($a^{\dagger}b+ab^{\dagger}$). While the former generates entanglement, and the latter cools the mechanical resonator to sustain the entanglement. In Fig. 2(b), the effect of the bath temperature on the optomechanical entanglement is further elucidated by fixing $\tilde{\Delta}_{a}=0.9\omega_{b}$ and $G=4\kappa_{a}$. It is shown that the optomechanical entanglement monotonically decreases as the bath temperature increases. The survival temperature for the optomechanical entanglement can be up to $\sim 190$ mK.
@@
Figure 3: Density plot of $E_{N}^{ab}$ vs the normalized frequency detuning $\Delta_{m}/\omega_{b}$ and the normalized magnon-photon coupling strength $g_{m}/G$ with (a) $\kappa_{m}=\kappa_{a}$ and (b) $\kappa_{m}=10\kappa_{m}$. (c) The optomechanical entanglement ($E_{N}^{ab}$), the magnon-photon entanglement ($E_{N}^{am}$), and the magnon-phonon entanglement ($E_{N}^{bm}$), vs the normalized magnon frequency detuning $\Delta_{m}/\omega_{b}$, where $g_{m}=G$ and $\kappa_{m}=\kappa_{a}$. (d) The magnon-photon-phonon entanglement ($\mathcal{R}_{\tau}^{abm}$) VS the normalized magnon frequency detuning $\Delta_{m}/\omega_{b}$ with different magnon-photon coupling strengths $g_{m}/G=0.2,0.4,1$, where $\kappa_{m}=\kappa_{a}$ is taken. Other parameters are the same as those in Fig. 2(a) except for $G=4\kappa_{a}$ and $\tilde{\Delta}_{a}=0.9\omega_{b}$.
@@ III.2 With magnons
To investigate the magnon-mediated optomechanical entanglement, we take the hybrid cavity-magnon optomechanics with a single sphere as an example, that is, we set $g_{1}=g_{m}\neq 0$ and $g_{2}=0$ in Fig. 1. Correspondingly, $\Delta_{1}=\Delta_{m}$ is specified. Within the framework of cavity magnonics, the parameters such as the magnon frequency, the magnon-photon coupling, and hte magnon decay rate are all tuned, leading to tunable optomechanical entanglement generated, as numerically investigated in Fig. 3. From Fig. 3(a), it is observed that the optomechanical entanglement decreases with increasing the magnon-photon coupling while keeping other system parameters fixed. Around $\Delta_{m}\approx\omega_{b}$, the optomechanical entanglement is optimal when $g_{m}$ is comparable to $G$ ($g_{m}\sim G$). This can also be evident in Fig. 3(b). Notably, a larger decay rate of the Kittle mode can significantly preserve the optomechanical entanglement, which can be given by comparing Figs. 3(a) and 3(b). Furthermore, Figure 3(c) elucidates that the optomechanical entanglement will be reduced when magnons are introduced. The reduction of the optomechanical entanglement is initially transformed into the magnon-photon entanglement (blue curve) and subsequently into the magnon-phonon entanglement (red curve). To simultaneously achieve these three entanglements, tuning the detuning of the cavity mode to be an optimal value is demanded. Additionally, we demonstrate that the magnon-photon-phonon tripartite entanglement ($\mathcal{R}_{\tau}^{abm}$) around $\Delta_{m}\approx-\omega_{b}$ can be anticipated, as depicted in Fig. 3(d).By increasing the magnon-photon coupling, this tripartite entanglement noticeably improves, indicating that the optomechanical entanglement is also partially transfered to the tripartite entanglement among the Kittle mode, the cavity mode, and the mechanical mode.
@@ IV.3 Noise effect
In Fig. 6(a), the effects of the magnon decay rate on both the magnon-magnon and the optomechanical entanglement are investigated, where $g_{m}=G$ and $\Delta_{1}=-\Delta_{2}=-\omega_{b}$ are fixed. It is observed that larger decay rates of the magnons do not promote the generation of magnon-magnon entanglement. Instead, they facilitate the production of optomechanical entanglement, consistent with the findings from Figs. 3(a) and 3(b). This not only reveals that the entanglement exchange between these two entanglement occurs, but also shows that the generated magnon-magnon entanglement can be be leveraged to restore the optomechanical entanglement by increasing the decay rate of the Kittle mode. Since the bath temperature is crucial for entanglement survival, we respectively plot $E_{N}^{m_{1}m_{2}}$ and $E_{N}^{ab}$ vs the bath temperature $T$ in Figs. 6(b) and 6(c) with different magnon-photon coupling strengths, where $\kappa_{m}=\kappa_{a}$ is fixed. When the magnon-photon coupling strength is weaker than the optomechanical coupling strength [see $g_{m}=0.5G$ in Fig. 6(b)], the optomechanical entanglement (the black curve) is stronger than the magnon-magnon entanglement (the red curve) within the survival temperature. But when the magnon-photon coupling strength is comparable with the optomechanical coupling strength [see $g_{m}=G$ Fig. 6(c)], the situation is reversed, that is, the magnon-magnon entanglement becomes much stronger than the optomechanical entanglement. Moreover, the survival temperature of the optomechanical entanglement is significantly shifted with varying the magnon-photon coupling. This indicates that the survival temperature of the magnon-magnon entanglement is more robust against the change of the magnon-photon coupling than that of the optomechanical entanglement.
@@ V.2 Magnon-magnon-phonon entanglement
Next, we go on studying the magnon-magnon-phonon entanglement in the proposed hybrid cavity-magnon optomechanics. Note that there are no direct magnon-magnon and magnon-phonon couplings. We observe that optimal entanglement occurs around $\Delta_{1}=\Delta_{2}\approx-0.5\omega_{b}$. This point coincides with the nose of the fox-pattern depicted in Fig. 8(a). A comparison with Fig. 7(a) reveals that the magnon-magnon-phonon entanglement is stronger than that of magnon-magnon-photon entanglement, which can be regarded as that the mechanical mode can be cooled by the optomechanical beam-splitter interaction, or alternatively, the noise effect on the magnon-magnon-phonon entanglement is weaker than the magnon-magnon-photon entanglement because of the indirect interaction among magnons and phonons. We also find that such a tripartite entanglement can be improved by enhancing the magnon-photon coupling [see Fig. 8(b)]. Upon examining the entanglement with respect to the decay rate of the Kittle mode, we plot the magnon-magnon-phonon entanglement ($\mathcal{R}_{\tau}^{bm_{1}m_{2}}$) vs the normalized decay rate $\kappa_{m}/\kappa_{a}$. It is shown that a weaker decay rate does not benefit the generation of magnon-magnon-phonon entanglement. By gradually increasing the decay rate, the magnon-magnon-phonon entanglement increases to its maximum and then decreases to zero. Comparing Figs. 7(c) and 8(c), we find that both the magnon-magnon-phonon entanglement and the magnon-magnon-photon entanglement have the similar resilience against the magnon decay rate. However, the magnon-magnon-phonon entanglement exhibits more robust againt the bath temperature thant that of the magnon-magnon-photon entanglement, as shown in the insets of Figs 7(c) and 8(c). In fact, the magnon-magnon-phonon entanglement is about two times than the magnon-magnon-photon entanglement within the survival temperature180similar-toabsent180$\sim$ 180∼ 180 mK.
@@
In summary, we propose to generate diverse bipartite and tripartite entanglement in hybrid cavity-magnon optomechanics. Unlike standard cavity optomechanics, our proposal not only enables magnon-mediated optomechanical entanglement but also allows for flexible magnon-magnon entanglement. It can even generate tripartite entanglement, including magnon-photon-phonon, magnon-magnon-photon, and magnon-magnon-phonon configurations. Moreover, optimal bipartite and tripartite entanglement can be achieved by tuning parameters. We further show that all entanglement can be enhanced via engineering the magnon-photon coupling, and are robust against the bath temperature within the survival temperature. In addition, we find that the optomechanical entanglement can be protected or restored by bad magnons with large decay rate, while other entanglement are significantly diminished. These findings suggest that our proposal offers a novel avenue to explore and control tunable macroscopic quantum effects in hybrid cavity-magnon optomechanics.
This work is supported by Zhejiang Provincial Natural Science Foundation of China under Grant No. LY24A040004.

### v4

@@ II.3 Covariance matrix
where $\mathcal{D}=\operatorname{diag}[\kappa_{1}(2\bar{n}_{1}+1),\kappa_{1}(2\bar{n}_{1}+1),\kappa_{2}(2\bar{n}_{2}+1),\kappa_{2}(2\bar{n}_{2}+1),\kappa_{a}(2\bar{n}_{a}+1),\kappa_{a}(2\bar{n}_{a}+1),0,\gamma_{b}(2\bar{n}_{b}+1)]$ is defined by $\langle\bar{n}_{l}(t)\bar{n}_{k}(t^{\prime})$ $\bar{n}_{k}(t^{\prime})\bar{n}_{l}(t)\rangle=2\mathcal{D}_{lk}\delta(t-t^{\prime})$.
@@ III Optomechanical entanglement
In this part, we study the magnon-mediated optomechanical entanglement. To be more clear, we first re-examine the optomechanical entanglement without magnons, i.e., $g_{1}=0$ and $g_{2}=0$. Then we investigate the magnon-mediated optomechanical entanglement by taking a single sphere for example. Without loss of generality, we assume $g_{1}=g_{m}\neq 0$ and $g_{2}=0$.
Table 1: Symbols used in numerical calculation
Symbol Meaning $\omega_{a}/2\pi=10~\mathrm{GHz}$ Cavity frequency $\omega_{b}/2\pi=10~\mathrm{MHz}$ Mechanical frequency $\omega_{1(2)}/2\pi=10~\mathrm{GHz}$ Magnon frequency $\kappa_{a}/2\pi=1~\mathrm{MHz}$ Cavity decay rate $\gamma_{b}/2\pi=100~\mathrm{Hz}$ Mechanical decay rate $\kappa_{1(2)}/2\pi=1~\mathrm{MHz}$ Magnon decay rate $\omega_{0}/2\pi=9.95~\mathrm{GHz}$ Frequency of the driving field $\Omega/2\pi=0.4~\mathrm{GHz}$ Amplitude of the driving field $T=20~\mathrm{mK}$ Bath temperature $\tilde{\Delta}_{m}=0.9\omega_{b}$ Cavity detuning $\Delta_{1(2)}=5\omega_{b}$ Magnon detuning
@@ III.1 Without magnons
Figure 2: (a) The optomechanical entanglement ($E_{N}^{ab}$) vs the dimensionless normalized cavity detuning $\tilde{\Delta}_{a}/\omega_{b}$ with different optomechanical coupling strengths $G/\kappa_{a}=1,3,5$ at the bath temperature $T=20$ mK, where $\tilde{\Delta}_{a}$ is the cavity detuning and $\omega_{b}$ is the frequency of the mechanical mode. (b) The LN $E_{N}^{ab}$ vs the bath temperature $T$ with the moderate optomechanical coupling strength $G/\kappa_{a}=4$ and the frequency detuning $\tilde{\Delta}_{a}=0.9\omega_{b}$. Other parameters are chosen as $\omega_{a}/2\pi=10$ GHz, $\omega_{b}/2\pi=10$ MHz, $\kappa_{a}/2\pi=1$ MHz, $\gamma_{b}/2\pi=100$ Hz.
@@
When no spheres are placed in the cavity (i.e., $g_{1}=g_{2}=0$), the proposed cavity-magnon optomechanics reduces to a standard optomechanical configuration comprising solely a cavity and a mechanical resonator. In this setup, optomechanical entanglement can be analyzed by plotting LN ($E_{N}^{ab}$) vs the normalized cavity detuning $\tilde{\Delta}_{a}/\omega_{b}$ with different linearized coupling strengths $G$ in Fig. 2(a). Here, we employ experimentally accessible parametersZhang et al. (2016); Bernier et al. (2017): $\omega_{a}/2\pi$ = 10 GHz, $\omega_{b}/2\pi$ = 10 MHz, $\kappa_{a}/2\pi=1$ MHz, $\gamma_{b}/2\pi=10^{2}$ Hz, and $T=20$ mK. Besides, other parameters such as $\omega_{0}=9.95$ GHz and $\Omega=0.4$ GHz are chosen. Evidently, the optomechanical entanglement increases first and then decreases with the cavity detuning $\tilde{\Delta}_{a}/\omega_{b}$. Around $\tilde{\Delta}_{a}\approx\omega_{b}$, the optimal optomechanical entanglement is observed. By gradually increasing the optomechanical coupling, the optomechanical entanglement can be significantly enhanced (as depicted by the three curves). The generated optomechanical entanglement arises from both the two-mode squeezing effect ($a^{\dagger}b^{\dagger}+ab$, where $q\propto b+b^{\dagger}$) and the beam-splitter effect ($a^{\dagger}b+ab^{\dagger}$). While the former generates entanglement, and the latter cools the mechanical resonator to sustain the entanglement. In Fig. 2(b), the effect of the bath temperature on the optomechanical entanglement is further elucidated by fixing $\tilde{\Delta}_{a}=0.9\omega_{b}$ and $G=4\kappa_{a}$. It is shown that the optomechanical entanglement monotonically decreases as the bath temperature increases. The survival temperature for the optomechanical entanglement can be up to $\sim 190$ mK.
@@
Figure 3: Density plot of $E_{N}^{ab}$ vs the normalized frequency detuning $\Delta_{m}/\omega_{b}$ and the normalized magnon-photon coupling strength $g_{m}/G$ with (a) $\kappa_{m}=\kappa_{a}$ and (b) $\kappa_{m}=10\kappa_{a}$. (c) The optomechanical entanglement ($E_{N}^{ab}$), the magnon-photon entanglement ($E_{N}^{am}$), and the magnon-phonon entanglement ($E_{N}^{bm}$), vs the normalized magnon frequency detuning $\Delta_{m}/\omega_{b}$, where $g_{m}=G$ and $\kappa_{m}=2\kappa_{a}$. (d) The optomechanical entanglement vs the normalized frequency detuning $\Delta_{m}/\omega_{b}$ with different values of $\kappa_{m}/\kappa_{a}=2,5,10$. The star means the optimal entanglement. (e) The magnon-photon-phonon entanglement ($\mathcal{R}_{\tau}^{abm}$) vs the normalized magnon frequency detuning $\Delta_{m}/\omega_{b}$ with different magnon-photon coupling strengths $g_{m}/G=0.2,0.6,1$, where $\kappa_{m}=\kappa_{a}$ is taken. Other parameters are the same as those in Fig. 2(a) except for $G=4\kappa_{a}$ and $\tilde{\Delta}_{a}=0.9\omega_{b}$.
@@ III.2 With magnons
To investigate the magnon mediated optomechanical entanglement, we take the hybrid cavity-magnon optomechanical system with a single YIG sphere as an example, that is, we set $g_{1}=g_{m}\neq 0$ and $g_{2}=0$ in Fig. 1, and accordingly specify $\Delta_{1}=\Delta_{m}$. Within the framework of cavity magnonics, parameters such as the magnon frequency, the magnon-photon coupling strength, and the magnon decay rate can be tuned, resulting in controllable optomechanical entanglement. The corresponding numerical results are shown in Fig. 3. Figure 3(a) shows how the optomechanical entanglement varies with the magnon-photon coupling strength $g_{m}$ while keeping other system parameters fixed. The entanglement gradually decreases as $g_{m}$ increases. Nevertheless, an optimal entanglement value appears when $g_{m}$ becomes comparable to the effective optomechanical coupling $G$, that is, $g_{m}\sim G$. This optimal behavior is further confirmed in Fig. 3(b), which demonstrates that the optomechanical entanglement reaches its maximum around the same parameter range. Moreover, a comparison between Figs. 3(a) and 3(b) reveals that a larger magnon decay rate can help preserve or even recover the optomechanical entanglement, highlighting the nontrivial role of the dissipative magnons in the hybrid system. The effect of the magnon decay rate is further examined in Fig. 3(c). It is shown that introducing magnons with a small decay rate leads to a noticeable reduction in the optomechanical entanglement. This reduction is accompanied by a redistribution of quantum correlations into magnon–photon (blue curve) and magnon-phonon (red curve) entanglements, particularly when the cavity detuning approaches $\Delta_{m}\approx-\omega_{b}$. Figure 3(d), which magnifies the pink dashed region in Fig. 3(c), shows that the optimal optomechanical entanglement occurs at $\Delta_{m}=0.64\omega_{b}$ when $\kappa_{m}=2\kappa_{a}$ (black curve). As the magnon decay rate increases ($\kappa_{m}=5\kappa_{a}$ and $10\kappa_{a}$), the optimal point gradually shifts toward $\Delta_{m}\approx\omega_{b}$, indicating that stronger magnon damping alters the resonance condition of the hybrid modes. Finally, Fig. 3(e) displays the magnon photon phonon tripartite entanglement $\mathcal{R}_{\tau}^{abm}$, which becomes pronounced around $\Delta_{m}\approx-\omega_{b}$. Increasing the magnon photon coupling significantly enhances this tripartite entanglement, implying that part of the optomechanical entanglement is converted into genuine three mode correlations among the Kittel mode, the cavity mode, and the mechanical mode.
@@ IV.3 Noise effect
In Fig. 6(a), the effects of the magnon decay rate on both the magnon-magnon and the optomechanical entanglement are investigated, where $g_{m}=G$ and $\Delta_{1}=-\Delta_{2}=-\omega_{b}$ are fixed. It is observed that larger decay rates of the magnons do not promote the generation of magnon-magnon entanglement. Instead, they facilitate the production of optomechanical entanglement, consistent with the findings from Figs. 3(a) and 3(b). This not only reveals that the entanglement exchange between these two entanglement occurs, but also shows that the generated magnon-magnon entanglement can be leveraged to restore the optomechanical entanglement by increasing the decay rate of the Kittle mode. Since the bath temperature is crucial for entanglement survival, we respectively plot $E_{N}^{m_{1}m_{2}}$ and $E_{N}^{ab}$ vs the bath temperature $T$ in Figs. 6(b) and 6(c) with different magnon-photon coupling strengths, where $\kappa_{m}=\kappa_{a}$ is fixed. When the magnon-photon coupling strength is weaker than the optomechanical coupling strength [see $g_{m}=0.5G$ in Fig. 6(b)], the optomechanical entanglement (the black curve) is stronger than the magnon-magnon entanglement (the red curve) within the survival temperature. But when the magnon-photon coupling strength is comparable with the optomechanical coupling strength [see $g_{m}=G$ Fig. 6(c)], the situation is reversed, that is, the magnon-magnon entanglement becomes much stronger than the optomechanical entanglement. Moreover, the survival temperature of the optomechanical entanglement is significantly shifted with varying the magnon-photon coupling. This indicates that the survival temperature of the magnon-magnon entanglement is more robust against the change of the magnon-photon coupling than that of the optomechanical entanglement.
@@ V.2 Magnon-magnon-phonon entanglement
Next, we investigate the magnon-magnon-phonon entanglement in the proposed hybrid cavity-magnon optomechanical system. It should be emphasized that there are no direct couplings between the two magnon modes or between the magnons and the phonon. Instead, the entanglement among them arises from indirect interactions mediated by the cavity field, which acts as an effective quantum bus Sarma et al. (2021). The optimal magnon-magnon-phonon entanglement is observed around $\Delta_{1}=\Delta_{2}\approx-0.5\omega_{b}$, corresponding to the tip of the fox-shaped pattern shown in Fig. 8(a). Compared with Fig. 7(a), the magnon-magnon-phonon entanglement is stronger than the magnon-magnon-photon entanglement. This can be attributed to two factors. First, the mechanical mode can be effectively cooled through the optomechanical beam-splitter interaction, which enhances its quantum correlations with the magnons Aspelmeyer et al. (2012); Bowen and Milburn (2015). Second, the noise effect on the magnon-magnon-phonon entanglement is weaker because the magnons and phonons interact indirectly, which helps to suppress decoherence in this subsystem Li et al. (2018).
Moreover, Fig. 8(b) indicates that the tripartite magnon-magnon-phonon entanglement can be further enhanced by increasing the magnon-photon coupling strength, consistent with previous studies of hybrid quantum systems Qiu et al. (2022); Wang et al. (2015). To explore the influence of the magnon decay rate, we plot the magnon-magnon-phonon entanglement ($\mathcal{R}_{\tau}^{bm_{1}m_{2}}$) as a function of the normalized decay rate $\kappa_{m}/\kappa_{a}$. The results show that a very small decay rate does not favor the formation of tripartite entanglement. As the decay rate increases, the entanglement first rises to a maximum and then gradually vanishes when losses dominate. Comparing Figs. 7(c) and 8(c), both the magnon-magnon-phonon and magnon-magnon-photon entanglements exhibit similar resilience to the magnon decay rate. However, the magnon-magnon-phonon entanglement demonstrates stronger robustness against the bath temperature, as shown in the insets of Figs. 7(c) and 8(c). Specifically, within the survival temperature range of approximately 180 mK, the magnitude of the magnon-magnon-phonon entanglement is nearly twice that of the magnon-magnon-photon entanglement. Note that the degree of tripartite entanglement achieved here remains weak. However, it could be further enhanced by introducing coherent feedback loop Amazioug et al. (2023), parametric amplifier Hussain et al. (2022), optical squeezing Zhang et al. (2022), combination of parametric and squeezing Wodedo et al. (2025), mechanical nonlinearity Wan et al. (2024), and three-level atoms Kussia et al. (2024). Incorporating these effects may provide an efficient route to strengthen multipartite quantum correlations, which will be an interesting topic for future investigation. To highlight the novelty of our scheme, a comparison with these existing schemes is provided in Table 2.
Table 2: Comparison of our work with others
Reference Method Mechanism $\mathbf{E_{N}^{m_{1}m_{2}}}$ $\boldsymbol{\mathcal{R}_{\tau}^{m_{1}m_{2}a(b)}} R$ef. Amazioug et al. (2023) Feedback loop Magnomechanics No No Ref. Hussain et al. (2022) Parametric amplifier Magnomechanics No No Ref. Zhang et al. (2022) Squeezing Magnomechanics No No Ref. Wodedo et al. (2025) Parametric + squeezing Optomechanics No No Ref. Wan et al. (2024) Mechanical nonlinearity Optomechanics No No Ref. Kussia et al. (2024) Three-level atoms Optomechanics No No Our work Magnon Optomechanics Yes Yes
@@
[…]
In summary, we propose to generate diverse bipartite and tripartite entanglement in hybrid cavity-magnon optomechanics. Unlike standard cavity optomechanics, our proposal not only enables magnon-mediated optomechanical entanglement but also allows for flexible magnon-magnon entanglement. It can even generate tripartite entanglement, including magnon-photon-phonon, magnon-magnon-photon, and magnon-magnon-phonon configurations. Moreover, optimal bipartite and tripartite entanglement can be achieved by tuning parameters. We further show that all entanglement can be enhanced via engineering the magnon-photon coupling, and are robust below the survival threshold temperature. In addition, we find that the optomechanical entanglement can be protected or restored by bad magnons with large decay rate, while other entanglement are significantly diminished. These findings suggest that our proposal offers a novel avenue to explore and control tunable macroscopic quantum effects in hybrid cavity-magnon optomechanics.
This work was supported by the Natural Science Foundation of Zhejiang Province (GrantNo. LY24A040004), Zhejiang Province Key R&D Program of China (Grant No. 2025C01028), and Shenzhen International Quantum Academy (Grant No. SIQA2024KFKT010).